### v6

# RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Interactive Environmental Learning in Physical Embodied Systems

Abstract
Embodied intelligence aims to enable robots to learn, reason, and generalize robustly across complex real-world environments. However, existing approaches often struggle with partial observability, fragmented spatial reasoning, and inefficient integration of heterogeneous memories, limiting their capacity for long-horizon adaptation. To address this, we introduce RoboMemory, a brain-inspired framework that unifies Spatial, Temporal, Episodic, and Semantic memory within a parallelized architecture for efficient long-horizon planning and interactive learning. Its core innovations are a dynamic spatial knowledge graph for scalable, consistent memory updates and a closed-loop planner with a critic module for adaptive decision-making. Extensive experiments on EmbodiedBench show that RoboMemory, instantiated with Qwen2.5-VL-72B-Ins, improves the average success rate by 26.5% over its strong baseline and even surpasses the closed-source SOTA, Claude-3.5-Sonnet. Real-world trials further confirm its capability for cumulative learning, with performance consistently improving over repeated tasks. Our results position RoboMemory as a scalable foundation for memory-augmented embodied agents, bridging insights from cognitive neuroscience with practical robotic autonomy.
Figure 1: RoboMemory adopts a brain-inspired architecture that maps neural components to agent modules, enabling long-term planning and interactive learning across diverse environments (real-world, Habitat, ALFRED) and robotic hardware.
**footnotetext: Equal contribution
I Introduction
Recent advances in Vision-Language Models (VLMs) [23, 6] have enabled their growing use in embodied tasks [33, 22]. VLM-based embodied agents can process multimodal inputs and generate high-level textual commands (e.g., “Pick up the cup”), which require translation via tool APIs to become executable robot actions. In contrast, Vision-Language-Action models (VLAs) [24, 8, 7, 13] produce low-level control signals directly but generally rely only on the latest observation. This limits their ability to perform long-horizon, multi-step tasks that require reasoning over task history. In summary, VLA models enable direct robot control but lack high-level planning capabilities, and VLM-based embodied agents support strategic planning but struggle with direct motor control. This highlights a key gap inherent in two distinct technical approaches to embodied intelligence.
To bridge this gap, recent work [50, 34, 38] proposes a “VLM planner + VLA executor” paradigm. Here, VLM-based embodied agents serve as high-level planners that decompose complex tasks (e.g., “make a coffee”) into executable sub-instructions (e.g., “grasp the cup”) that VLAs can complete. Although this paradigm improves performance on multi-step tasks, prior work suffers from two key limitations in real-world settings. First, real-world tasks (e.g., kitchen operations) require navigating across multiple locations to gather objects and tools, but the environment remains only partially observable at any time due to robots’ limited field of view and dynamic occlusions. This necessitates a planner with robust spatial awareness and long-term memory to maintain a consistent spatial awareness across viewpoints. However, most VLM-based agents rely on chat-style context windows (e.g., logging instruction – feedback pairs [49]), which lack mechanisms for maintaining an overview of the environment’s spatial layout. Consequently, agents cannot reliably track object locations or recognize previously visited states. Second, pretrained VLMs are rarely trained on embodied planning trajectories, especially long-horizon, spatially grounded ones. So VLM-based agents often struggle to generalize to real-world settings [47]. To overcome these challenges, VLM-based planners must support interactive environmental learning — the ability to acquire, integrate, and retrieve spatial, episodic, and semantic knowledge during task execution, thereby enabling adaptation through experience. In summary, the current VLM–VLA paradigm still lacks (1) a robust spatial and long-term memory mechanism to maintain a consistent view of the environment, and (2) an effective way to acquire and leverage embodied planning experience for generalization.
To address these challenges, we propose RoboMemory, a brain-inspired memory framework designed to bridge the gap between biological cognitive mechanisms and robotic embodied intelligence. Drawing on foundational cognitive theories, we recognize that human memory is inherently organized into a tiered architecture rather than a flat buffer: sensory memory captures transient perceptual inputs [37], short-term working memory handles immediate information manipulation [4, 5], and long-term memory provides persistent storage for episodic events and semantic facts [3, 43]. Neuroscientific research further reveals that specialized brain regions underpin these functions: the thalamus integrates multimodal sensory inputs, the hippocampus consolidates spatial and episodic experiences into long-term storage [29], the prefrontal cortex regulates high-level planning based on retrieval, and the cerebellum coordinates low-level motor execution. Mimicking this biological hierarchy to ensure layered information processing, our recipe for this generalist memory system consists of four key stages (Figure 1): 1) Information Preprocessing (Thalamus-inspired), which integrates multimodal sensory inputs for downstream processing; 2) Comprehensive Embodied Memory (Hippocampus-inspired), which organizes experiential and spatial knowledge through a three-tier structure (sensory, short-term, and long-term). Crucially, we design four parallel-update modules (Spatial, Temporal, Episodic, and Semantic) to minimize latency while ensuring data consistency; 3) Closed-Loop Planning (Prefrontal Cortex-inspired), where the planner utilizes retrieved memory to generate high-level action sequences; 4) Low-level Execution (Cerebellum-inspired), which translates plans into precise robot actions using VLA models and SLAM-based navigation.
While recent frameworks [39, 18, 45, 1, 16, 54, 10] have integrated Retrieval-Augmented Generation (RAG) to enhance planning, they suffer from a critical limitation: most are tailored for simulated environments or rely on flat, text-based retrieval that fails to capture the spatial complexity of the physical world. These approaches often lack a hierarchical structure, leading to inefficiencies in long-term knowledge accumulation. Moreover, existing multi-module systems [38, 1] typically incur significant inference latency due to sequential memory updates. In contrast, our framework leverages a parallel-update paradigm and a hierarchical organization, enabling the agent to maintain a dynamic, spatially-aware worldview in real-time, which is essential for robust interaction in dynamic environments.
We evaluate RoboMemory on EmbodiedBench, a challenging long-horizon planning benchmark [47]. Using Qwen2.5-VL-72B as the base model, RoboMemory achieves state-of-the-art results, improving average success rates by 26.5% over the baseline and outperforming the closed-source Claude3.5-Sonnet [2]. In real-world deployments, our system demonstrates the capacity for continuous improvement: by executing diverse tasks sequentially without memory resets, the agent effectively learns from environmental familiarization.
Our main contributions are three-fold:
  • We propose RoboMemory, a brain-inspired, unified embodied memory system that integrates four parallel update modules into a single framework. It enables efficient, comprehensive memory operations and coherent knowledge integration, which are critical for interactive environmental learning in real-world embodied scenarios.
  • We design a retrieval-based incremental update algorithm for the real-time evolution of Spatial Knowledge Graphs (KGs). This algorithm bridges the gap between static knowledge representation and dynamic embodied interaction, establishing KGs as a practical and scalable modality for Spatial Memory. By retrieving relevant subgraphs, detecting local inconsistencies, and merging new observations, our method ensures efficient maintenance and overcomes the scalability bottlenecks that previously hindered KG-based approaches.
  • RoboMemory supports interactive environmental learning in real-world environments. It enables sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements, demonstrating practical long-term autonomous learning in physical scenarios.
Figure 2: (a) Left: The loop where the Planner, Critic, and Comprehensive Embodied Memory interact to adjust plans based on real-time visual inputs. Colored text denotes the execution status of actions (success/rejected/replanned). (b) Right: Spatial memory maintains a relevance/similarity-updated KG, and Semantic/Episodic memory manages a Vector DB with analogous logic. Besides, Temporal memory is implemented as a linear FIFO buffer that stores step-wise summaries generated by the Step Summarizer.
II Related Work
II-A VLM/LLM-based Agentic Frameworks in Embodied Tasks
The rapid advancement of VLMs/LLMs has led to diverse agent frameworks in embodied environments [49, 36, 26]. Embodied tasks involve partial observability and long-horizon planning, requiring memory systems to retain context. Some use time-ordered context buffers for short-term memory (due to VLMs/LLMs’ limited long-context processing) [49, 32]; others adopt experience buffers as long-term semantic memory [16, 35]. For long-duration tasks, skill libraries serve as procedural memory, with agents accumulating skills via interaction [45, 39]. However, in real-world settings, the low-level executor may fail to complete the task, making it challenging to construct a reusable, code-based skill library. So, explicit procedural memory still needs to be improved in real-world settings. Moreover, Recent efforts integrate diverse memories [51, 39, 1] but focus on virtual/GUI environments, leaving real-world multi-modal memory support for long-term planning under-explored.
II-B Spatial Memory
Spatial information is critical for embodied agents. Prior work has explored three main representations, each with inherent limitations that motivate our high-level abstractions: 2D Semantic Maps encode environments as pixel grids [51]. As the semantic maps are flattened, they discard 3D relationships (e.g., “above/beneath”) and precise geometry, limiting expressiveness for manipulation tasks requiring fine-grained spatial reasoning. 3D Point Clouds preserve accurate 3D information [52]. However, a 3D point cloud only encodes low-level spatial information. We need to put in extra effort to help agents understand 3D Point Clouds. Semantic Scene Graphs [38, 28, 19, 48, 20] offering a compact, high-level representation directly amenable to planning. However, existing methods suffer from two key drawbacks: (1) Most previous methods adopt rigid tree-like structures, which are easy to update with new information, but cannot model non-hierarchical or cyclic spatial configurations. (2) Some previous methods try to use general graphs to store various spatial information. [20] However, their design often lacks efficient dynamic update mechanisms—adding or modifying relations typically requires expensive full-graph recomputation, creating a scalability bottleneck for real-time embodied interaction.
For clarity, we summarize the differences among different memory systems in the previous work. The comparison is shown in Table III in the appendix. Additionally, we provide more ablation studies in Appendix -A.
III RoboMemory
RoboMemory is a hierarchical embodied agent system that equips robots with three core memory capabilities: historical interaction logs, dynamically updated spatial layouts, and accumulated task knowledge. As illustrated in Figure 2, in each iteration, RoboMemory follows a process of “Perception – Memory – Retrieval – Planning – Execution” process, ensuring that the agent continuously calibrates its memory and behavior in dynamic environments.
First, the information preprocessor converts multimodal sensor inputs into a textual summary of the current scene, which serves as the primary input to the Comprehensive Embodied Memory. Next, the Comprehensive Embodied Memory updates its internal representations, including action histories, object locations, and experiential knowledge. After updating the information, the memory system retrieves contextually relevant entries to inform the Closed-Loop Planning Module. Then, leveraging this contextual memory, the Closed-Loop Planning Module generates high-level, text-based action instructions. Finally, these commands are dispatched to low-level executors, which directly control the robot and complete the instructions. The execution process is demonstrated in Algorithm 1 in the appendix.
III-A Information Preprocessor
At each time step $t$, RoboMemory receives a visual observation $\mathcal{O}_{t}$: an RGB frame (in simulation) or a short video clip (on physical robots), representing the agent’s observations. Since raw visual data is unsuitable for direct use in memory construction and retrieval, RoboMemory first employs an information preprocessor to convert multimodal observations into textual representations, thereby providing a semantic interface for subsequent memory and planning modules. The information preprocessor executes two Vision-Language Models (VLMs) in parallel: (1) Step summarizer $\mathcal{S}$: It transforms $\mathcal{O}_{t}$ into a concise textual description $s_{t}$ of the just-executed action. The string $s_{t}$ is stored in the system’s working memory. (2) Query generator $\mathcal{Q}$: It derives a list of queries $q_{t}=[q_{t}^{(1)},q_{t}^{(2)},\dots,q_{t}^{(N)}]$ from the same observation $\mathcal{O}_{t}$. Each query $q^{(i)}_{t}$ is a natural language-based query. These queries are used to query information from the memory system that may be useful.
Together, $\mathcal{S}$ and $\mathcal{Q}$ provide a swift, text-based interface between raw sensory data and provide basic information in each iteration for RoboMemory’s Comprehensive Embodied Memory System.
III-B Comprehensive Embodied Memory
To address the long-term memory limitations in current embodied agent frameworks, we propose the Comprehensive Embodied Memory System. This system consists of multiple memory modules. We denote the memory system containing $L$ distinct modules as $M_{t}=[M_{t}^{(1)},M_{t}^{(2)},...,M_{t}^{(L)}]$, where $M_{t}$ represents the memory stored at step $t$, and $M_{t}^{(l)}$ denotes the $l$-th memory module. Generally, the update and retrieval process at iteration $t$ of a general memory is shown below:
$M_{t}=\mathcal{U}(M_{t-1},s_{t})$ (1)
$r_{t}=\mathcal{R}(M_{t},q_{t})$ (2)
First, we update memory modules with memory update algorithm $\mathcal{U}$, where we update the previous memory $M_{t-1}$ using the latest summarization $s_{t}$. Then, with updated $M_{t}$, we use memory retrieve algorithm $\mathcal{R}$ to retrieve the information that is useful for the planning module. In $\mathcal{R}$, we use queries $q_{t}$ to query $M_{t}$, yielding retrieval results from each module: $r_{t}=[r_{t}^{(1)},r_{t}^{(2)},...,r_{t}^{(L)}]$. These results are then passed to the planning module, which helps it plan future movements. However, sequentially updating and retrieving from $L$ modules would be a slow process. Therefore, we parallelize these steps across all modules, which significantly enhance the efficiency of the memory system.
The memory system consists of four distinct modules ($L=4$): Temporal Memory, Spatial Memory, Semantic Memory, and Episodic Memory. For efficiency, all memory modules are updated and retrieved in parallel. Thus, even with multiple modules, the system remains highly efficient. Functionally, inspired by cognitive psychology [27], our modules handle memory at different levels. In cognitive psychology, memory is divided into Sensory Memory, Short-term Memory, and Long-term Memory. Mirroring this hierarchy, our modules are organized as follows. First of all, the step summarizer $\mathcal{S}$ summarizes the agent’s interactions with the environment at each iteration. It acts as Sensory Memory. Secondly, Temporal Memory and Spatial Memory function as short-term memory. These two memories will be updated at each iteration. They are designed to store the information of sensory memory in every iteration. For Temporal Memory, we record the agent’s action history sequentially, while for Spatial Memory, we dynamically record the spatial relationships between different objects in the environment based on Sensory Memory. These memories can provide a relatively long and detailed history of the current task for the Closed-Loop Planner to make a future plan. Thirdly, Semantic Memory and Episodic Memory serve as Long-term Memory. They are updated only when meaningful information arises (e.g., after task completion). These memories store highly abstract knowledge, not limited to the current task, but synthesized from past experiences. This knowledge—factual, event-based, and experiential—improves the agent’s future task performance. It is the source of RoboMemory’s interactive learning capability. We now detail each module.
Temporal Memory. In the Temporal Memory, we record interactions between the robot and the environment (i.e., Sensory Memory) of each iteration sequentially. This information can provide the embodied agent with simple awareness of “What I have done”. For such temporally sequential memory, a simple structure is sufficient: a sequential buffer with automatic summarization triggered when the record sequence reaches its capacity. In a specific design, temporal memory can store up to $N$ interaction summaries, each generated by an information preprocessor. When the buffer is full, we compress the oldest $N$ steps into a single summarized entry using a VLM, which is then reinserted at the front of the buffer, ensuring continuous context retention without unbounded growth. However, the information from previous memories will gradually be lost as we summarize it multiple times. For retrieval, we provide all existing memory in text to downstream modules.
Spatial Memory. The spatial memory is designed to dynamically record the high-level spatial relationships of different entities in the environment. However, current spatial memory approaches often rely on RGB-D cameras to reconstruct 3D point clouds [51, 9]. These representations are too detailed for high-level planning in embodied agents. For example, precise geometric relationships (e.g., exact distances between objects) are unnecessary.
To address these problems, we use a dynamic KG to store high-level spatial information: objects and positions in the environment become vertices of the KG, and spatial relations between objects or positions are encoded as edges. The KG focuses on high-level spatial relations (e.g., “cup on table”, “key left of drawer”). By these settings, spatial KG focuses on semantically meaningful, task-relevant relations. This spatial information enhances the agent’s spatial reasoning capability in dynamic environments.
However, as related work shows, most KG construction algorithms are designed for static long content. This does not meet the demand of using KG as spatial memory for the agent. The KG needs to update efficiently in response to new information. To address this issue, we introduce a retrieval-driven, incremental KG update algorithm that maintains a locally modifiable, globally consistent, and dynamically adaptive spatial memory. As illustrated in the right panel of Figure 2, the update process proceeds in four steps: (1) retrieves the most relevant sub-KG around new observations. (2) Injects new relations from the current observation by a VLM-based Relation Retriever. (3) Detects and resolves conflicts between newly extracted relations and existing ones (e.g., “cup on table” vs. “cup in drawer”) using a VLM-based resolver, which decides whether to add, delete, or modify edges. (4) Merges back and prunes isolated vertices. Moreover, our retrieval-based incremental update algorithm is accompanied by provable efficiency guarantees. For a KG with $n$ vertices and maximum degree $D$, the number of vertices processed per update is bounded by $O(D^{K})$, where $K$ is the retrieval hop distance (see Appendix -D for analysis). Further implementation details are provided in Appendix -C1.
Semantic Memory. In cognitive psychology, semantic memory stores time-independent facts. These facts are stable, update slowly, and require long-term retention. In RoboMemory, semantic memory records task-relevant experiences and environmental knowledge during execution. This information can help RoboMemory adapt to new environments or tasks. This information is highly abstract and does not need to be updated frequently. In RoboMemory, semantic memory updates when new information is encountered during execution, for example, after completing a subtask or encountering important information. To store and update memory efficiently, we design a memory management system based on a vector database (DB). In the vector DB, each experience/fact is described in natural language (denoted as a memory item). Each memory item is converted into a semantic vector for querying. For dynamic updates, we adapt a basic method from prior work [12]. However, we adopt it into the embodied environment. As shown in the bottom-right of Figure 2, the semantic memory update algorithm involves two VLM-based modules. Firstly, the Long-term Memory Creator generates new memory items based on short-term memory. We retrieve the top-$S$ most similar existing memory items from existing memory via cosine similarity. A VLM-based updater then compares new and existing items to decide whether to: add the new item, update an existing item, remove an outdated item, or perform Noop (if redundant). Since updates only involve a maximum of $S$ previous memory items and the update process is parallelized across all memory modules, this update method ensures that semantic memory remains efficient even as the database grows. For retrieval, we use the same process as a traditional vector DB. We use queries to extract top-$N$ relevant information for downstream modules.
Episodic Memory. In cognitive psychology, Episodic Memory is another important part of long-term memory. It can store task-specific execution summaries (i.e., “autobiographical” records of past attempts). In RoboMemory, the Episodic Memory module is responsible for recording every interaction trajectory it has gone through, including the sequence of actions the robot did and the feedback from the environment. The trajectory information can help to improve the planning ability of RoboMemory. For example, if a trajectory for completing a similar task is stored in episodic memory, it can guide the agent in completing the current task. As the agent only needs to follow the successful trajectory in the memory, it can reduce hallucinations or errors in the VLM planner. Technically, Episodic Memory shares the same storage and VLM-driven vector DB update mechanism as Semantic Memory, ensuring consistent architectural design.
III-C Closed-Loop Planning Module
The Closed-Loop Planning Module integrates information about the current task provided by the Spatial-Temporal Memory, Semantic and Episodic information recorded in long-term memory, and current observations to perform action planning. Each action is planned and passed on to the low-level executor for execution.
To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism [25], which consists of the planner and the critic module which is powerd by same VLM model with different prompts. We denote the planner module as $\mathcal{P}$, while the critic module as $\mathcal{C}$. For each planning step, the planner generates a long-term plan consisting of multiple steps. However, due to the dynamics of embodied environments, the action sequence in the long-term plan may become outdated during the execution of the plan. Thus, before executing each step, we use the Critic model to evaluate whether the proposed action in this step remains appropriate under the latest environment. If not, the planner will re-plan based on the latest information. The demonstration of this process is shown in Figure 2.
However, our experiments reveal that the original Planner-Critic mechanism may suffer from infinite loops. In the original mechanism, the first step of the action sequence output by the Planner is evaluated by the Critic before execution, which can lead to an infinite loop: if the Critic always demands replanning, no action will ever be executed. To address this, we modify the Planner-Critic mechanism so that the first step is not evaluated by the Critic. This ensures that even if the Critic persistently demands replanning, the RoboMemory will still execute actions. The detailed algorithm is shown in Algorithm 1 in the appendix.
III-D Low-Level Executor
The RoboMemory is a two-layer hierarchical agent framework that accomplishes longer-term tasks in the real world. The upper layer is responsible only for high-level planning, while the Low-level Executor carries out the actions planned by the upper layer in the real environment.
We employ a LoRA-finetuned VLA model, $\pi_{0}$ [21, 8], to generate manipulation actions, and a SLAM-based navigation model for locomotion. The low-level executor then translates high-level actions planned by RoboMemory into concrete arm and chassis movements in the real world.
TABLE I: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels (Base/Long) on EB-ALFRED and EB-Habitat benchmarks. Values are reported in percentages (%).
Method Type EB-ALFRED EB-Habitat
Average Base Long Base Long
SR GC SR GC SR GC SR GC SR GC
Single VLM-Agents
GPT-4o Closed-source 67.0 74.9 64.0 74.0 54.0 62.5 86.0 90.7 64.0 72.2
GPT-4o-mini 30.5 40.9 34.0 47.8 0.0 17.0 74.0 77.5 14.0 21.3
Claude-3.7-Sonnet 68.5 - 68.0 - 70.0 - 90.0 - 46.0 -
Claude-3.5-Sonnet 69.5 71.8 72.0 72.0 52.0 54.5 96.0 97.5 58.0 63.3
Gemini-1.5-Pro 68.0 73.3 70.0 74.3 58.0 65.0 92.0 92.5 52.0 61.2
Gemini-2.0-flash 57.0 61.5 62.0 65.7 58.0 62.0 82.0 82.0 26.0 36.2
Llama-3.2-90B-Vision-Ins Open-source 40.5 46.6 38.0 43.7 16.0 24.0 94.0 94.5 14.0 24.3
InternVL2.5-78B 47.0 52.9 38.0 42.3 42.0 49.0 80.0 82.0 28.0 38.2
InternVL2.5-38B 37.5 42.6 36.0 37.3 26.0 36.5 60.0 61.5 28.0 35.0
InternVL3-78B 49.5 - 38.0 - 36.0 - 84.0 - 40.0 -
Qwen2.5-VL-72B-Ins 44.0 - 50.0 - 34.0 - 74.0 - 18.0 -
VLM-Agent Frameworks
Voyager (Qwen2.5-VL-72B-Ins) Baselines 46.5 66.4 56.0 73.2 32.0 54.2 76.0 87.0 22.0 51.0
Reflexion (Qwen2.5-VL-72B-Ins) 38.3 51.1 48.0 54.0 10.0 33.0 80.0 84.2 15.0 33.0
Cradle (Qwen2.5-VL-72B-Ins) 44.5 57.0 54.0 67.9 32.0 41.0 62.0 67.0 30.0 52.1
RoboOS (Qwen2.5-VL-72B-Ins) 25.5 33.0 32.0 38.4 12.0 17.6 38.0 47.8 20.0 28.2
RoboOS (RoboBrain2-32B) 20.0 25.7 32.0 37.2 8.0 13.2 28.0 34.8 12.0 17.4
RoboMemory (Qwen2.5-VL-72B-Ins) Ours 70.5 79.7 68.0 75.5 66.0 81.3 86.0 88.0 62.0 74.0
IV Experiments
IV-A Benchmarks
To evaluate RoboMemory’s task planning ability, we select a subset of the EmbodiedBench EB-ALFRED and EB-Habitat benchmark [47]. We select the Base and Long subsets because they aim to test the agent’s planning ability. The Base and Long subsets of the two benchmarks comprise 200 tasks for complex embodied tasks. The EB-ALFRED and EB-Habitat benchmarks provide a visually grounded operational setting that closely mimics real-world conditions (see Appendix -E for environment details), enabling direct comparison with established baselines. We set temperature = 0 to avoid randomness during the experiment, aligning with the EmbodiedBench setting. Each task is executed once. Moreover, we set up an environment to test the interactive environmental learning ability of RoboMemory in the real world.
IV-B Baselines and Metrics
To facilitate comparisons, we consider two types of baselines. First, we choose the advanced closed-source and open-source VLMs as a single agent. We compare their performance with RoboMemory. For closed source VLMs, we choose GPT-4o and GPT-4o-mini [31, 23], Claude3.5-Sonnet and Claude-3.7-Sonnet [2], Gemini-1.5-Pro and Gemini-2.0-flash [41, 15]. For open source VLMs, we choose Llama-3.2-90B-Vision-Ins [30], InternVL-2.5-78B/28B [11], InternVL-3-72B [55], and Qwen2.5-VL-72B-Ins [6]. Secondly, we choose three agent frameworks: (1) Reflexion [35], which introduces a simple long-term memory and a self-reflection module. Reflexion uses the self-reflection module to summarize experiences as long-term memory, thereby enhancing the model’s capabilities. (2) Voyager [45], which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning. (3) Cradle [39], which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks. (4) RoboOS [38], which proposes an embodied agent framework that consists of a Hierarchy Scene-Graph based Spatial Memory.
In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins [42] with temperature set as 0. For the RoboOS framework, we test it on RoboBrain2-32B [40], where the RoboBrain2-32B model is designed for the RoboOS framework.
The Qwen2.5-VL-72b-Ins represents a high-performing open-source alternative. Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks [46]. We use the Qwen3-Embedding model [53] to create embedding vectors for RAGs in RoboMemory. For the Low-level Executor, since EB-ALFRED provides high-level action APIs, we use the Low-level executor provided by EmbodiedBench instead of the VLA-based method.
We define two evaluation metrics to assess the performance: (1) Success Rate (SR), which is the ratio of completed tasks to the total number of tasks in each difficulty level. This metric reflects the agent’s ability to complete tasks across randomly generated scenarios. (2) Goal Condition Success Rate (GC), which is the ratio of intermediate conditions achieved to the maximum possible score in each scenario. An GC of 100% indicates that the task is completed in the given scenario. These two metrics can be computed as:
$SR=\mathbb{E}_{x\in\mathcal{X}}\left[\mathds{1}_{SCN_{x}=GCN_{x}}\right]$ (3)
$GC=\mathbb{E}_{x\in\mathcal{X}}\left[\frac{SCN_{x}}{GCN_{x}}\right]$ (4)
In above formulas, $\mathcal{X}$ denotes the test subset, and $x$ represents a test task. The success condition number ($SCN_{x}$) refers to the number of conditions the agent has accomplished, while the global condition number ($GCN_{x}$) indicates the total number of conditions required for task completion. The task is considered successful if $SCN_{x}=GCN_{x}$.
IV-C Main Results
Figure 3: Efficiency improvement of Comprehensive Embodied Memory System.
As shown in Table I, our model achieves significant improvements over both single VLM agents and Agent frameworks on the EB-ALFRED and EB-Habitat. Compared to the SOTA Single VLM-Agent model, Claude3.5-Sonnet, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 1% and GC by 7.9%. This demonstrates RoboMemory’s superiority over single VLM-Agents, proving that an Agent framework with open-source models can outperform closed-source SOTA models. Furthermore, when tested against other VLM-Agent frameworks, RoboMemory also shows substantial gains. This is because, unlike other agent frameworks, RoboMemory’s brain-like memory system provides embodied models with more accurate and persistent contextual information. Additionally, the Planner-Critic mechanism provides a closed-loop planning ability, which helps the RoboMemory gain better performance in long-term tasks. Because the RoboMemory can detect and try to overcome possible failures. And it is more robust when encountering unexpected situations.
IV-D Efficiency Analysis
To evaluate the efficiency of the Comprehensive Embodied Memory module, we tested RoboMemory with executing 10 long-horizon tasks, each comprising approximately 50 steps. All testing is under the same hardware and software settings. We exclusively measured the wall-clock time consumed by memory update and retrieval operations. We analyzed the scaling behavior of memory update latency across three distinct configurations: (1) fully parallel update and retrieval across all memory modules; (2) sequential update of each memory module without parallelism; and (3) update of only the most fundamental memory component: the Temporal Memory. Results are presented in Figure 3. As shown, our parallel update strategy enables updating a multi-module memory system with latency comparable to that of updating a single base memory module. This demonstrates the critical efficiency gains afforded by parallelization across the memory architecture.
TABLE II: Ablation study of RoboMemory’s Success Rate (SR) on EB-ALFERD subset.
Method (EB-ALFRED) Avg. Base Long
RoboMemory 67% 68% 66%
- w/o critic 55 % 60 % 50%
- w/o spatial memory 47 % 52 % 42 %
- w/o episodic memory 62% 68% 56%
- w/o semantic memory 58% 66% 50%
- w/o long-term memory 57% 66% 48%
Figure 4: Visualization of the experimental environment.
Figure 5: Real-world experiment results. “Qwen” denotes Qwen2.5-VL-72B-Ins; “RoboBrain” denotes RoboBrain2.0-32B.
IV-E Ablation Studies
We used the full Base and Long Subset from EB-ALFRED to validate RoboMemory’s effectiveness. We removed each component systematically and observed performance changes across task categories. We use the success rate as our metric. Results are shown in Table II.
Long-term Memory: Adding long-term memory significantly improves RoboMemory’s success rate. The experiment shows that it enables interactive environmental learning while attempting to complete tasks. Semantic memory learns the properties of low-level skills, such as in what circumstances an action may fail. The episodic memory records all task attempts (successful/failed), providing valuable experience at the task level and giving insight into how to complete a task successfully. This helps the RoboMemory predict the outcomes of actions and avoid ineffective attempts. This ability indicates that the RoboMemory has an interactive environmental learning capability. Additionally, we do ablations for semantic memory and episodic memory solely. The result shows that those memory modules significantly improve the capability of the performance of RoboMemory.
Spatial Memory: Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability. Our novel dynamic KG update algorithm enables KG-based spatial memory in dynamic environments. This spatial reasoning helps RoboMemory handle partially observable embodied settings.
Critic Module: Table II shows performance without the critic module (55% vs 67% with full system). This drop highlights how the critic’s closed-loop planning adapts to dynamic environments. It helps RoboMemory recover from failures faster and handle unexpected situations better.
IV-F Real-World Robot Deployment
To evaluate RoboMemory’s interactive environmental learning capability in the real world, we designed a kitchen environment inspired by EB-ALFRED and EB-Habitat. The scene contains 5 navigable points, 8 interactive objects, and over 10 non-interactive (but potentially distracting) items. The environment is shown in Figure 5. In the real world, we use interactive environmental video recordings captured during action execution (rather than static snapshots taken after action completion) as RoboMemory’s input. This provides a more temporally coherent perception. We created three task categories (5 tasks each). Tasks are matched to EB-ALFRED’s Base subset (avg. oracle: 10–20 steps), though actual executions often exceed 20 steps due to search and error recovery. Due to search and error recovery, the robot often exceeds 20 steps per task. Additional hardware experiment details are in Appendix -E.
To test the interactive environmental learning ability of RoboMemory, we ran each task twice without clearing long-term memory between attempts. Meanwhile, we compared RoboMemory against previous SOTA on real-world experiments RoboOS as baselines. The success rates for first and second attempts and different settings of RoboOS are shown in Figure 5.
The second attempt showed significantly higher success rates. This proves RoboMemory’s long-term memory effectively guides subsequent tasks in real embodied environments. Key observations include: (1) Closed-loop error recovery: RoboMemory retries failed actions when possible, even if the low-level executor (VLA model) fails. (2) Spatial reasoning: RoboMemory remembers object locations and spatial relationships using its memory. (3) Interactive environmental learning: RoboMemory analyzes failure causes reasonably. These analyses guide future decisions. Detailed examples demonstrating these capabilities and further discussions are provided in Appendix -F.
Moreover, we observe a significant drop in task success rates when deploying the agent with the Low-level Executor in real-world environments. This performance degradation primarily stems from the executor’s inherent limitations: (1) The VLA model exhibits unreliable instruction-following capabilities, frequently failing during grasping actions or selecting incorrect objects; (2) Pre-trained VLM models demonstrate inadequate video understanding capability. They are struggle to interpret dynamic visual information such as action failures or state changes. These limitations collectively contribute to the reduced performance compared to simulated environments.
V Conclusion
This paper propose RoboMemory, a brain-inspired multi-memory framework, facilitating long-horizon planning and interactive environmental learning in real-world embodied systems by addressing key challenges such as memory latency, task correlation capture, and planning loops. Experiments on EmbodiedBench demonstrate that RoboMemory outperforms state-of-the-art closed-source VLMs and agent frameworks, with ablation studies confirming the critical roles of the Critic module and spatial/long-term memory. Real-world deployment further validates its interactive learning capability through improved success rates in repeated tasks. Despite limitations arising from reasoning errors and executor dependence, RoboMemory provides a foundation for generalizable, memory-augmented agents, with future work aimed at refining reasoning and enhancing execution robustness.